40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Mar 20 12:42:01 2015, Jenne, Update, IOO, Waking up the IFO 
    Reply  Fri Mar 20 13:29:33 2015, Koji, Update, IOO, Waking up the IFO 
    Reply  Fri Mar 20 16:44:49 2015, ericq, Update, IOO, Waking up the IFO 
       Reply  Fri Mar 20 23:37:46 2015, Jenne, Update, SUS, Waking up the IFO 
          Reply  Sat Mar 21 05:19:49 2015, Jenne, Update, LSC, IFO awake carm_cm_up_zip.sh.gz
          Reply  Sun Mar 22 13:53:44 2015, rana, Update, SUS, ITMX alignment jumps 
Message ID: 11150     Entry time: Fri Mar 20 12:42:01 2015     Reply to this: 11151   11152
Author: Jenne 
Type: Update 
Category: IOO 
Subject: Waking up the IFO 

I've done a few things to start waking up the IFO after it's week of conference-vacation.

PMC trans was at 0.679, aligned the input to the PMC, now it's up at 0.786.

MC transmission was very low, mostly from low PMC transmission.  Anyhow, MC locked, WFS relieved so that it will re-acquire faster.

Many of the optics had drifted away. AS port had no fringing, and almost every optic was far away from it's driftmon set val.  While putting the optics back to their driftmon spots, I noticed that some of the cds.servos had incorrect gain.  Previously, I had just been using the ETMX servo, which had the correct gain, but the ITMs needed smaller gain, and some of the optics needed the gain to be negative rather than positive.  So, now the script ..../scripts/SUS/DRIFT_MON/MoveOpticToMatchDriftMon.py has individually defined gains for the cds.servo. 

Next up (after lunch) will be locking an aligning the arms.  I still don't have MICH fringing at the AS port, so I suspect that the ASS will move some of the optics somewhat significantly (perhaps the input tip tilts, which I don't have DRIFT_MON for?)

ELOG V3.1.3-